FInd the demand functions for each of the ff.:
1. U = x1^a x2^b
2.U = min{ax1, bx2}
3. U = ax1 + bx2)
1
Expert's answer
2016-12-04T04:29:10-0500
1. U = x1^a x2^b. The demand functions are:x1=am/(p1(a+b)), x2=bm/(p1(a+b)), where m is income, p1, p2 - are prices of the goods 1 and 2 in the set.
2.U = min{ax1, bx2}. The demand functions are:x1=bm/(bp1+ap2), x2=am/(bp1+ap2).

3. U = ax1 + bx2). The demand functions are: x1=m/p1, x2=0 if a/p1>b/p2 or x1=0, x2=m/p2 if b/p2>a/p1.